# Diversity and stylization of the contemporary user-generated visual arts in the complexity-entropy plane

### v2

@@ 1 Introduction
Second, while theoretical studies of creativity in the domain of art can be traced back to earlier eras, contemporary efforts to quantitatively assess and measure relevant characteristics have only recently gained momentum [6, 7]. Through the development of computational and robust methods for analyzing visual arts and the accumulation of large-scale digital art images, scientists have characterized and uncovered various interesting features within a broad range of artworks. Computational studies of visual art have provided both synchronic and diachronic insights into zeitgeist and evolution of visual arts. Previous studies of visual art styles have been advanced with statistical data of spatial parameters, including chromatic properties [8], fractals [9], wavelets [10], compression ensemble approach [11], and concepts of entropy [12, 13, 14].
@@
2) We use two types of similarity measures - cosine similarity and Jaccard similarity, on two multiple image representation space – image embeddings through a pre-trained Residual Network (ResNet) architecture [19] and Scale-invariant Feature Transform (SIFT) features [20], to measure the image diversity/degree of dissimilarity in a specific C-H region. The low- (e.g., visual elements such as lines, contours, height, edges, angles, dots, colors, etc.) and high-level image features (e.g., themes of shapes and objects comprised of low-level features) extracted from the two methods aggregately encompass art style-level explicability of images [21].
3) In addition, we use the autoregressive moving-average (ARMA) model to statistically examine a temporal relationship between the average C-H positions of artworks within a given period and the average dissimilarities of their multi-level image features. Finally, we investigate the diversity of image representations that can be expressed in different areas of the C-H space and unveil an empirical condition for the emergence of styles that are not only novel in the C-H plane, but also characterized by greater stylistic diversity.
@@ 2.1 Stylization of curated visual arts in the C-H plane: DeviantArt and Bēhance (2010-2020)
To account for the stylization of contemporary visual artworks mainly distributed via online communication channels of information, we investigate DeviantArt and Bēhance, serving as massive online visual art platforms involving various creative fields. Since there are huge collections of artworks in both platforms, we specifically focus on representative and quasi-canonical subsets of daily promoted user-generated visual artworks (“Daily Deviation” and “Best of Bēhance”) curated by those platforms. This allowed us to process manageable and significant amounts of data [22, 23]. The image set used for complexity-entropy (C-H) measurement is described in Table SI 1.
@@
Figure 1: Group-level temporal transitions of visual art styles in the C-H plane. (a) Contemporary user-generated visual art styles of DeviantArt and Bēhance (2010 - 2020) compared to the main divisions of art historical periods mapped in the C-H plane. The result reveals that the average C-H values of art historical styles have been shifted towards the higher C and lower H area over time. (b, c) Yearly C-H trajectory of contemporary user-generated visual arts from DeviantArt’s Daily Deviations and Bēhance’s the Best of Bēhance (refer to Data (Materials)). Each C-H plane is composed of multiple elliptical areas, with each ellipse representing 95% confidence interval bounds from the yearly average C-H values of the intragroup visual artworks. The multiple CI ellipses and the yearly C-H trajectories connecting them altogether describe yearly transitions of visual art styles within each platform. Also, insets are included to show cumulative number of samples over the given time frame.
@@
Fig. 1 demonstrates that both platforms’ C-H trajectories are tended towards the upper left C-H region compared to the previous periods, although their average position shows a slight difference (refer to Table SI 2 for more information regarding the difference in the overall positions of the two platforms in the C-H plane).
@@
Also, a closer look on the C-H values of DeviantArt and Bēhance reveals that the averaged position of the two platforms have varied significantly over time. We cross-validated the temporal difference in C-H values on each platform by evaluating the predictive accuracies of our dataset’s C-H values using four machine learning algorithms. All the classification models predict the curation year of paintings with probabilities (DeviantArt: $\sim$15%, Bēhance: $\sim$13.7%) significantly higher than the chance level (Fig. SI 1).
@@
The C-H trajectories of temporal transitions of DeviantArt and Bēhance visual art styles over a decade extend that of the transition between early 20th century and 1970s (e.g., from painterly/optic to linear/haptic). Observing the beginning and ending years (Fig. 1b, c), the C of DeviantArt’s visual art increased by 0.002 (from 0.123 to 0.125) while their H decreased by 0.005 (from 0.84 to 0.835). In the case of Bēhance, there was a rise in C by 0.01 (from 0.12 to 0.13) and a decline in H by 0.11 (from 0.81 to 0.7) over the same period. The group-level temporal transitions of visual art styles among the two platforms reveal indirect but eventual tendencies towards styles with higher C and lower H. Consequently, the chronological C-H trajectories of the contemporary user-generated visual arts of both platforms from 2010 to 2020 indicate that their stylizations altered in a macroscopically similar manner, despite the differences in direction details within each platform.
@@
Meanwhile, the yearly average C-H values of Bēhance are notably distinct from those of DeviantArt and the major art historical periods. We explore which sub-visual art fields substantially influenced the C-H positions of Bēhance over time, by separately examining the temporal C-H movements of five creative fields that account for the greatest proportion of samples in Bēhance dataset. We observe the yearly average C-H values of samples from Illustration, Graphic Design, Character Design, Animation, and Fine Arts fields in Bēhance. Fig. SI 2 shows that all the five fields’ yearly average C-H values are specifically directed towards a similar C-H area (C: 0.1 - 0.3, H: 0.65 - 0.85) through a decade of movements. Overall, the C-H movements of the fields reveal a tendency for homogeneous transitions of visual art styles in Bēhance during the time period.
@@
Extending Sigaki et al.’s analysis on the evolution of art styles allows us to discover the group-level temporal stylizations of quasi-canonical visual artworks in DeviantArt and Bēhance (2010 – 2020) aligning with that of the modern art identified by Sigaki et al. Overall, the C-H movements of the platforms are characterized by an increase in high-level visual structure [25] of artworks incorporating geometrical object-oriented patterns; the emergence and taking up of an artistic style with clearer and simpler visual elements.
@@
Through structural notions of creativity and stylization, it is possible to infer the strengthened stylization of the contemporary user-generated visual arts towards a more linear/haptic region in the C-H plane. One possible interpretation is the “bijectively homomorphic” (i.e., isomorphic) [26] stylization of the artworks. In other words, random and diverse visual artistic representations of artworks interact and homogenize while linearly transforming underlying characteristics of art-makings over time. The significance of individual artists’ innovation being proportional to the extent of its influence on others [27], fuels a competitive and mimetic process under uncertainty, paradoxically resulting in a cycle of imitation as a necessary step toward creativity [28]. Also, in any given institution, the number of “insiders” on the curatorial committee is fewer than expected, resulting in “a durable and recognizable pattern of aesthetic choices [29]”. In response to the inherent “symbolic uncertainty” in the platform of visual arts, individual artists may be subject to “mimetic isomorphism,” or mimicking others’ distinguished artistry as the most cost-effective strategy for gaining insider acceptance [30]. One could argue that the constantly quickened dissemination of information and images, influenced and expected by visual content structures on the platforms, has accelerated the homogenization of contemporary user-generated visual art styles.
@@ 2.2 Spatiotemporal relationships between image diversity and the C-H space
Figure 2: Image representations in the complexity-entropy (C-H) plane and pipelines of image similarity measures through different image representation spaces. (a) Three reputable exemplary masterpieces chosen and localized in the C-H plane according to the given embedding parameters (dx = dy = 2), to aid viewers’ understandings of the image representations the plane encompasses. (b) A schematic diagram of image similarity computation pipeline through a neural network. We measure pairwise cosine similarities of a 200-dimensional embedding vector including both low- and high-level feature maps extracted from individual visual artworks. (c) A schematic diagram of image similarity computation pipeline through SIFT descriptor. We measure pairwise Jaccard similarities of SIFT descriptor matching a 128-dimensional vector of low-level features extracted from individual visual artworks. In addition to the image representations the C-H plane encompasses, we adopt and use two different types of image processing methods to obtain multi-level image features and measure their similarities: ResNet architecture and SIFT algorithm. Through our approach, we investigate characteristics of image representations in the C-H plane with image similarity measures that aggregate both low- and high-level features. The implemented images (1 – 3.) are “Kompozicija II by Piet Mondrian, 1929 (WikiArt),” “Rhythm by Robert Delaunay, 1912 (WikiArt),” and “The starry night by Vincent van Gogh, 1889 (WikiArt).” All the used image samples are available in the public domain.
We first examine the relationship between the average C-H positions of user-generated visual arts from DeviantArt and Bēhance over the specified timeframe (2010-2020) and their intragroup image diversity. In doing so, we apply the ARMA model to statistically regress changes in average values from the two image similarity measures on the average values of the complexity C and entropy H over the given years.
@@
Semirobust Standard errors in parentheses; *** p<0.001, ** p<0.01, * p<0.05 * + p<0.1
Table 1: ARMA predicting average similarity among visual arts in DeviantArt and Bēhance
The ARMA model is used separately for subsampled images from DeviantArt and Bēhance datatset. The results (Table 1) indicate that C and H values of a particular year are associated with an increase in diversity (over the previous year); the C and H values can affect the increase in diversity, despite the decrease in average absolute diversity over time. Here, we note that as the analysis predicts a movement of the average value of diversity for a given year with its average value of C and H, the area with large C and H values in the entire C-H space does not represent the area of high diversity. Meanwhile, the variance of H is added as an independent variable and controlled on the one hand, confirming that H variance also significantly has a positive correlation with diversity in all cases.
@@
A distribution of given years’ images in the C-H space is likely to include a wider variety of styles along with an increase in image diversity within the given year group upon the following conditions. First, when an average movement (direction of increases in C and H) occurs toward the upper right C-H region, which is the optimal direction of the upper left (similar, unexplored, sparse) and lower right (diverse, explored, dense) C-H region. Additionally, when the distribution (large H variance) is made on the entropy H axis.
We now partially examine the robustness of the ARMA results through patterns of the local stylistic diversity visualized over binned areas of a partitioned C-H region (C: 0 - 0.31, H: 0.5 – 1; see Fig. 3). Overall, the sample count plots (Fig. 3a, d) show that the most densely populated area is to the lower right C-H region with given embedding parameters, dx = dy = 2. This confirms that there is likely to be a strong correlation between image diversity and the highly concentrated C-H region, where previous art-historical styles with great image diversity are positioned. Hence, the more artworks where entropy H is greater, the greater image diversity there is likely to be.
@@
As observed in Fig. 1a, our results uncover the C-H movements of quasi-canonical visual artworks in contemporary online platforms, which ultimately trended to the upper-left C-H region. The C-H movements begin and propel away from the highly dense and stylistically diverse (low C and high H), and move towards sparser and stylistically homogenic areas (high C and low H), forming a process of a particular stylization. Such a stylization (towards higher C and lower H) is consistent with the broader movements in Western art history [13], which ranged from classical to modern times. However, the region with high C and low H has the lowest diversity (i.e., high image similarity) of image representations in artworks. Therefore, even though propelling towards the area is a new artistic endeavor in terms of C and H, there are few artworks that could be expressed with the given C-H values.
Based on our analyses, we suggest conditions for the groups of visual artworks to have high intragroup diversity, and shed light on an evolutionary process in which the intragroup diversity gives rise to styles. It is confirmed that when the average C-H movement of the groups are balanced towards higher C and higher H rather than the extremes in the C-H space, the stylistic diversity of the groups show a significant tendency to increase more compared to the previous period. Our findings also suggest that a visual artistic stylization process occurs as a result of the cumulative diversity of individualized artworks contributing as original mutations over time, to the transformative process of the prior narratives of styles, and emergence of a new style.

### v3

# Investigating the diversity and stylization of contemporary user generated visual arts in the complexity entropy plane

@@ 1 Introduction
Second, while theoretical studies of creativity in the domain of art can be traced back to earlier eras, contemporary efforts to quantitatively assess and measure relevant characteristics have only recently gained momentum [6, 7]. Through the development of computational and robust methods for analyzing visual arts and the accumulation of large-scale digital art images, scientists have characterized and uncovered various interesting features within a broad range of artworks. Computational studies of visual art have provided both synchronic and diachronic insights into the zeitgeist and evolution of visual arts. Previous studies of visual art styles have been advanced with statistical data of spatial parameters, including chromatic properties [8], fractals [9], wavelets [10], the compression ensemble approach [11], and concepts of entropy [12, 13, 14].
@@
2) We use two types of similarity measures—cosine similarity and Jaccard similarity—on two multiple image representation spaces—image embeddings through a pre-trained Residual Network (ResNet) architecture [19] and Scale-invariant Feature Transform (SIFT) features [20], to measure the image diversity/degree of dissimilarity in a specific C-H region. The low- (e.g., visual elements such as lines, contours, height, edges, angles, dots, colors, etc.) and high-level image features (e.g., themes of shapes and objects comprised of low-level features) extracted from the two methods aggregately encompass art style-level explicability of images [21].
In addition, we use the autoregressive moving-average (ARMA) model to statistically examine a temporal relationship between the average C-H positions of artworks within a given period and the average dissimilarities of their multi-level image features. Finally, we investigate the diversity of image representations that can be expressed in different areas of the C-H space and unveil an empirical condition for the emergence of styles that are not only novel in the C-H plane, but also characterized by greater stylistic diversity.
@@ 2.1 Stylization of curated visual arts in the C-H plane: DeviantArt and Bēhance (2010-2020)
To account for the stylization of contemporary visual artworks mainly distributed via online communication channels of information, we investigate DeviantArt and Bēhance, serving as massive online visual art platforms involving various creative fields. Since there are huge collections of artworks on both platforms, we specifically focus on representative and quasi-canonical subsets of daily promoted user-generated visual artworks (“Daily Deviation” and “Best of Bēhance”) curated by those platforms. This allowed us to process manageable and significant amounts of data [22, 23]. The image set used for complexity-entropy (C-H) measurement is described in Table SI 1.
@@
Figure 1: Group-level temporal transitions of visual art styles in the complexity-entropy (C-H) plane. (a) Contemporary user-generated visual art styles of DeviantArt and Bēhance (2010 - 2020) compared to the main divisions of art historical periods mapped in the C-H plane. Images with distinct forms and a limited number of ordinal patterns typically exhibit low H values and high C values. Conversely, disordered images comprising more random or interrelated components generally display high H values and low C values. The result reveals that the average C-H values of art historical styles have been shifted towards the higher C and lower H areas over time. (b, c) Yearly C-H trajectory of contemporary user-generated visual arts from DeviantArt’s Daily Deviations and Bēhance’s the Best of Bēhance (refer to Data (Materials)). Each C-H plane is composed of multiple elliptical areas, with each ellipse representing 95% confidence interval bounds from the yearly average C-H values of the intragroup visual artworks. The multiple CI ellipses and the yearly C-H trajectories connecting them altogether describe yearly transitions of visual art styles within each platform. Also, insets are included to show the cumulative number of samples over the given time frame.
@@
Fig. 1 demonstrates that both platforms’ C-H trajectories tend towards the upper-left C-H region compared to the previous periods, although their average position shows a slight difference (refer to Table SI 2 for more information regarding the difference in the overall positions of the two platforms in the C-H plane).
@@
Also, a closer look at the C-H values of DeviantArt and Bēhance reveals that the averaged positions of the two platforms have varied significantly over time. We cross-validated the temporal difference in C-H values on each platform by evaluating the predictive accuracies of our dataset’s C-H values using four machine learning algorithms. All the classification models predict the curation year of paintings with probabilities (DeviantArt: $\sim$15%, Bēhance: $\sim$13.7%) significantly higher than the chance level (Fig. SI 1).
@@
The C-H trajectories of temporal transitions of DeviantArt and Bēhance visual art styles over a decade extend that of the transition between early 20th century and 1970s (e.g., from painterly/optic to linear/haptic). Observing the beginning and ending years (Fig. 1b, c), the C of DeviantArt’s visual art increased by 0.002 (from 0.123 to 0.125), while their H decreased by 0.005 (from 0.84 to 0.835). In the case of Bēhance, there was a rise in C by 0.01 (from 0.12 to 0.13) and a decline in H by 0.11 (from 0.81 to 0.7) over the same period. The group-level temporal transitions of visual art styles among the two platforms reveal indirect but eventual tendencies towards styles with higher C and lower H. Consequently, the chronological C-H trajectories of the contemporary user-generated visual arts of both platforms from 2010 to 2020 indicate that their stylizations altered in a macroscopically similar manner, despite the differences in direction details within each platform.
@@
Meanwhile, the yearly average C-H values of Bēhance are notably distinct from those of DeviantArt and the major art historical periods. We explore which sub-visual art fields substantially influenced the C-H positions of Bēhance over time by separately examining the temporal C-H movements of five creative fields that account for the greatest proportion of samples in the Bēhance dataset. We observe the yearly average C-H values of samples from Illustration, Graphic Design, Character Design, Animation, and Fine Arts fields in Bēhance. Fig. SI 2 shows that all five fields’ yearly average C-H values are specifically directed towards a similar C-H area (C: 0.1 - 0.3, H: 0.65 - 0.85) through a decade of movements. Overall, the C-H movements of the fields reveal a tendency for homogeneous transitions of visual art styles in Bēhance during the time period.
@@
Extending Sigaki et al.’s analysis on the evolution of art styles allows us to discover the group-level temporal stylizations of quasi-canonical visual artworks in DeviantArt and Bēhance (2010 – 2020) aligning with that of the modern art identified by Sigaki et al. Overall, the C-H movements of the platforms are characterized by an increase in the high-level [25] visual structure of artworks incorporating geometrical object-oriented patterns; the emergence and taking up of an artistic style with clearer and simpler visual elements.
@@
Through structural notions of creativity and stylization, it is possible to infer the strengthened stylization of the contemporary user-generated visual arts towards a more linear/haptic region in the C-H plane. This trend can be interpreted as a process where artworks exhibit what Birkhoff called a “bijectively homomorphic” (structurally similar) [26] stylization. In simpler terms, while the visual representations of artworks remain diverse and random at an individual level, they interact and gradually align with one another, leading to a homogenization of styles. The significance of individual artists’ innovation being proportional to the extent of its influence on others [27], fuels a competitive and mimetic process under uncertainty, paradoxically resulting in a cycle of imitation as a necessary step toward creativity [28]. Also, in any given institution, the number of “insiders” on the curatorial committee is fewer than expected, resulting in “a durable and recognizable pattern of aesthetic choices [29]”. In response to the inherent “symbolic uncertainty” in the platform of visual arts, individual artists may be subject to “mimetic isomorphism,” or mimicking others’ distinguished artistry as the most cost-effective strategy for gaining insider acceptance [30].
The convergence of two distinct groups of images within the C-H framework over a decade exemplifies mimetic isomorphism, where shared environmental factors (e.g., technological trends, aesthetics, or cultural shifts) drive contextual adaptations rather than mere replication. This phenomenon highlights how stylistic and structural similarities can emerge across distinct entities through imitative and adaptive processes in complex creative domains. The rapid dissemination of information and images, driven by platform-based visual content structures, might have accelerated the homogenization of contemporary user-generated visual art styles.
@@ 2.2 Spatiotemporal relationships between image diversity and the C-H space
Figure 2: Image representations in the complexity-entropy (C-H) plane and pipelines of image similarity measures through different image representation spaces. (a) Three reputable exemplary masterpieces chosen and localized in the C-H plane according to the given embedding parameters ($d_{x}=d_{y}=2$), to aid viewers’ understandings of the image representations the plane encompasses. (b) A schematic diagram of an image similarity computation pipeline through a neural network. We measure pairwise cosine similarities of a 200-dimensional embedding vector including both low- and high-level feature maps extracted from individual visual artworks. (c) A schematic diagram of the image similarity computation pipeline through the SIFT descriptor. We measure pairwise Jaccard similarities of SIFT descriptor matching a 128-dimensional vector of low-level features extracted from individual visual artworks. In addition to the image representations the C-H plane encompasses, we adopt and use two different types of image processing methods to obtain multi-level image features and measure their similarities: ResNet architecture and SIFT algorithm. Through our approach, we investigate characteristics of image representations in the C-H plane with image similarity measures that aggregate both low- and high-level features. The implemented images (1 – 3.) are “Composition 2” by Piet Mondrian, 1929 (Public Domain, retrieved from WikiArt, https://www.wikiart.org/en/piet-mondrian/composition-2), “Rhythm” by Robert Delaunay, 1912 (Public Domain, retrieved from WikiArt, https://www.wikiart.org/en/robert-delaunay/rhythm-1), and “The Starry Night” by Vincent van Gogh, 1889 (Public Domain, retrieved from WikiArt, https://www.wikiart.org/en/vincent-van-gogh/the-starry-night-1889). All the used image samples are available in the public domain.
We first examine the relationship between the average C-H positions of user-generated visual arts from DeviantArt and Bēhance over the specified timeframe (2010-2020) and their intragroup image diversity. To account for trend components in the time-series data, we use an ARMA model to regress the average values from the dyadic similarities among the artworks on the average values of the complexity C and entropy H over the given years.
The ARMA model is applied separately to subsampled images from the DeviantArt and Bēhance datasets. The results (Table 1) indicate that the C and H values of a particular year are associated with a decrease in similarity, suggesting that the C and H values can affect the increase in diversity, despite the decrease in average absolute diversity over time. Here, we note that as the analysis predicts a movement of the average value of diversity for a given year with its average value of C and H, the area with large C and H values in the entire C-H space does not represent the area of high diversity. Meanwhile, the variance of H is added as an independent variable and controlled on the one hand, confirming that H variance also significantly has a positive correlation with diversity in all cases.
@@
Semirobust Standard errors in parentheses: *** p<0.001, ** p<0.01, * p<0.05 * + p<0.1
Table 1: ARMA predicting average similarity among visual arts in DeviantArt and Bēhance. Refer to Methods; Statistical regression of visual artwork image similarities for detailed description.
@@
A distribution of given years’ images in the C-H space is likely to include a wider variety of styles along with an increase in image diversity within the given year group upon the following conditions. First, when an average movement (direction of increases in C and H) occurs toward the upper-right C-H region, which is the optimal direction of the upper-left (similar, unexplored, sparse) and lower-right (diverse, explored, dense) C-H region. Additionally, when the distribution (large H variance) is made on the entropy H axis.
We now partially examine the robustness of the ARMA results through patterns of the local stylistic diversity visualized over binned areas of a partitioned C-H region (C: 0 - 0.31, H: 0.5 - 1; see Fig. 3). Overall, the sample count plots (Fig. 3a, d) show that the most densely populated area is to the lower-right C-H region with given embedding parameters, $d_{x}=d_{y}=2$. This confirms that there is likely to be a strong correlation between image diversity and the highly concentrated C-H region, where previous art-historical styles with great image diversity are positioned. Hence, the more artworks where entropy H is greater, the greater image diversity there is likely to be.
@@
As observed in Fig. 1a, our results uncover the C-H movements of quasi-canonical visual artworks in contemporary online platforms, which ultimately trended to the upper-left C-H region. The C-H movements begin and propel away from the highly dense and stylistically diverse (low C and high H) and move towards sparser and stylistically homogenic areas (high C and low H), forming a process of a particular stylization. Such a stylization (towards higher C and lower H) is consistent with the broader movements in Western art history [14], which ranged from classical to modern times. However, the region with high C and low H has the lowest diversity (i.e., high image similarity) of image representations in artworks. Therefore, even though propelling towards the area is a new artistic endeavor in terms of C and H, there are few artworks that could be expressed with the given C-H values.
Based on our analyses, we suggest conditions for the groups of visual artworks to have high intragroup diversity and shed light on an evolutionary process in which the intragroup diversity gives rise to styles. It is confirmed that when the average C-H movement of the groups is balanced towards higher C and higher H rather than the extremes in the C-H space, the stylistic diversity of the groups shows a significant tendency to increase more compared to the previous period. Our findings also suggest that a visual artistic stylization process occurs as a result of the cumulative diversity of individualized artworks contributing as original mutations over time, to the transformative process of the prior narratives of styles and the emergence of a new style.